24.03 Paper on Rationalist Approaches

Matthew Gray

May 9, 1996

Introduction

In both anthropology and philosophy is a meaningful question to ask how one defines rationality, and given that, how one can identify rational beliefs. It seems clear that in some cases, individuals who seem irrational at first will, upon closer examination, turn out to be quite rational beings. How is it that one can explain the initial perception of irrationality however? If apparent irrationality is explainable such that it is in fact rational, does rationality maintain any meaning?

Dan Sperber discusses a number of explanations for what might be percieved as irrational behavior. Each of these will be described and discussed and examined as to how they might be applied in anthropological cases, such as the Azande beliefs about Witchcraft. Alternative explanations and views of rationality will be discussed as well.

Rationalizing the Irrational

Sperber discusses a number of approaches to rationalizing apparently irrational beliefs. First are the cases of Intellectualism and Symbolism, which are relatively straightforward. Sperber goes on to propose a more complex Relativist approach. Finally Sperber proposes a Rationalist approach which is likely the most subtle and complex of the explanations.

Intellectualism

The Intellectualist approach is relatively straightforward. Its essential allegation is that an individual may have an erroenous, but rational, belief. To another individual who recognizes the belief as erroneous, the belief in this obviously false belief (to them) seems irrational.

This explanation, though simple, is fairly potent at explaining away large amounts of apparently irrational belief. The typical example of how this works is the case of the belief that the Earth is flat. Certainly, with the limited amount of information that many people throughout history have had, this is a reasonable, and wholly rational belief. If these beings had all the information, they would clearly see that the Earth was, in fact, round.

Sperber discounts this approach as particularly useful by citing cases where cultures would fabricate complex systems which he alleges only complicate the nature of the world, and are therefore not “rational”, at least as one can reasonably explain with the Intellectualist approach.

Symbolism

The Symbolist approach is also quite simple, but only seems to explain a very particular type of percieved rationality. Essentially the symbolist view claims that apparently irrational behavior is explainable as not in fact the literal, irrational belief that it seems to be, but instead is a metaphorical belief that is not nearly so irrational when properly interpreted.

Sperber points out that this explanation is only usable in a very limited number of sitations. Attempts have been made to use this explanation in a more general sense, but typically are not terribly useful.

Relativism

Sperber then arrives at a relativistic approach in which he claims that “fundamental concepts ... are culturally determined”. As a result of this it seems that much apparent irrationality is in fact a catastrophic failure of translation.

That is, when someone seems to be stating something irrational, they are in fact stating something so radically different from what you interpreted them to be saying, that it is not irrational. This approach, like many forms of extreme relativism, has the problem of almost eliminating the possibility of something being identifiable as irrational. If there is a regular and expected catastrophic failure of translation, it is very difficult to label any beliefs as irrational with any degree of certainty.

Rationalism

Sperber further proposes a view he calls Rationalism which relies on a distinction between believing that a statement is true, and believing that the semantic content of that statment is true. Sperber calls the former case “representational beliefs” and goes on to say that a seemingly irrational being or culture could in fact be holding wholly rational representational beliefs.

This experiences some of the same problems as the Relativist view, in that it obviate the need for a concept of rationality at all. If a set of beliefs that are irrational based on their semantic value but can be considered rational if the beliefs are in fact only the “representational” beliefs in the statements themselves, then it becomes difficult to identify anything as genuinely irrational.

Sperber does however admit that there can be certain beliefs which are not “representational” and are “factual”. The only distinction he provides is that a “factual belief” is one where there is “awareness of (what to the subject is) a fact”. This seems to say that if the subject believes that his beliefs are facts then they are not representational. This seems a weak distinction.

Without a objective way of identifying a belief as either factual or representational, the approach seems to fall apart. Either some representations are a priori propositional representations representing factual beliefs, or it becomes necassary to ask what somebody believes about their beliefs in order to determine whether any given belief is factual or representational. This circle repeats and makes the entire line of reasoning implausible.

On the other hand, it would seem to be consistent the deny the existence of any factual beliefs (though Sperber would certainly not do this), and claim that all beliefs are complex representational beliefs.

An Anthropological Case: The Azande

In looking at a case like the Azande belief in witchcraft, one needs to identify the reasons why it seemed irrational in the first place. With the Azande, they held beliefs that were counter to much of the evidence they had, and it became necessary to make their beliefs more complex in order for it to continue to be consistent. An observer might claim that the fact that the Azande persist in believing in witchcraft, despite overwhelming evidence to the contrary, is irrational.

Of the approaches to explaining apparently irrational behavior discussed above, the Intellectualist approach seems best suited for the case of the Azande. Their beliefs in the existence of witches do seem quite rational, they simply happen to be in error.

It seems clear that the Symbolist approach does not apply in this case, and the Relativist approach seems wrong as well. While it is possible that there is extreme mistranslation of concepts going on in communication with the Azande, it genuinely appears that fairly coherent and accurate communication is going on.

The Rationalist approach does not seem to work well in this case either, as the Azande would certainly (to them) believe that their beliefs about witches are factual, and not at all committed to any particular representation.

Alternative Explanations and the Motiviation

for Explanation

Before trying to identify further explanations for explaining apparently irrational behavior, it makes sense to determine the motivation in doing so. If the goal is to provide a sufficiently holistic explanation such that “irrationality” is no longer a meaningful concept, one might as well solve the problem at the semantic level and explain away apparently irrational behavior by defining irrationality as something which cannot be. Therefore, if we believe something is irrational, we are mistaken, and it is in fact rational by some other explanation. This previous argument is, of course, an Intellectualist one.

Perhaps the closest explanation to what people actually think, is to take a modified version of the four approaches presented here. In addition to claiming that rationality requires consistency, it seems that most poeple’s definition of “rationality” also requires the acceptance of certain a priori tenets.

These basic, fundamental beliefs could be minimalistic (weak) or extremely thorough (strong). These a priori tenets could be taken to be the “factual beliefs” described by Sperber. Sperber, however, does not want to admit any a priori truths by which to judge rationality, so doing so would be a departure from Sperber’s Rationalism.

Equally, the Relativist and Intellectualist approaches can be enhanced by this added stipulation. While a “pure” Intellectualist might claim that someone could be rational (but mistaken) to deny their own existence, for example. Realistically though, this is likely an indication of irrationality. A Relativist gains in much the same way, but must first admit that generally there is not total failure of translation, but only partial failure.

These “weak” and “strong” forms of rationality allow one to span the entire spectrum of views on rationality. “Weak” rationality in its most extreme form, borders on defining away irrationality. “Strong” rationality, in its extreme form, nearly requires rationality to be defined as “those who think just like me”. Clearly both of these extremes are poor, but most people would seem to define it somewhere in between, providing a more useful meaning.

Conclusion

Each of the proposed explanations for apparent irrationality runs into some problems when taken to absurd cases. Intellectualism seems the most directly consistent, but Sperber is apparently dissatisfied with it’s ability to explain away certain cases of apparent irrationality. Symbolism isn’t terribly problematic, but isn’t terribly applicable, either. Relativism is a case that works well, but in an extreme form, denies the existance of irrationality.

Sperber’s Rationalism, particularly in combination with possibility of requiring certain “factual beliefs”, is powerful. In Sperber’s formulation, however, the distinction between “factual” beliefs and “representational” beliefs is defined in terms of a groups beliefs about their beliefs. This yields problems for his argument.